OpenStudy (anonymous):

Christopher orders a 3 topping pizza for $15.25, and a 5 topping pizza for $17.75. Write and solve system of linear equations to find the price of a plain cheese pizza ( no toppings) and the cost of each topping.

OpenStudy (shane_b):

x+3y=15.25 x+5y=17.75 Where x = the price of a plain cheese pizza. Can you solve for x and y from here?

OpenStudy (anonymous):

y is one topping on the pizza x is just cheese with no toppings since the toppings are extra you want to add however many toppings to the price of the cheese pizza, assuming they are the same price. the result should equal the two prices of the cheese pizzas, shane already gave the equations, so solve for x and y with a linear equation methodology

OpenStudy (anonymous):

you get 17.75 for the answer

OpenStudy (shane_b):

No you don't.

OpenStudy (anonymous):

yes even you put that

OpenStudy (shane_b):

No I didn't. I posted the equations to solve...the answer is not 17.75.

OpenStudy (shane_b):

@michealaludwick: Do you know how to solve this or do you need help?

OpenStudy (anonymous):

I have no idea! Please help!

OpenStudy (shane_b):

Ok, you have two equations based on what @jonnyrules96 and I posted above: x+3y=15.25 x+5y=17.75 There are different ways to solve this...here's one: Solve for one of the variables in one of the equations. I'll use the first equation and solve for x: x+3y=15.25 x=15.25-3y Now that I have x in terms in y, I can plug that equation back into the other equation in place of x: (15.25-3y)+5y=17.75 Solving for y you get 1.25...which is the cost per topping. Now that you know the cost of each topping, you can use any one of the equations to solve for x (the price of the cheese pizza without extra toppings): x+3(1.25)=15.25 Solving for x you get a price of 11.50 for just a plain cheese pizza.

OpenStudy (anonymous):

in all linear equations you just need to first solve for one variable and then use that variable and solve for the other ones. but you need more then one equation and the same amount of equations as variables

OpenStudy (anonymous):

thank you guys so much !

OpenStudy (shane_b):

Good luck :)